OpenStudy (anonymous):

Determine the dimension of a rectangular solid(with a square base )with maximum volume if its surface area is 337.5 square centimeter.

OpenStudy (anonymous):

maximum volume means height of solid = infinity :P

hartnn (hartnn):

are u aware of the concept of derivatives to find maximum values?

OpenStudy (anonymous):

i know how to derive but how to derive in maximum values definitely no

hartnn (hartnn):

ok, volume of rect solid with square base of side x and height y is x^2y did u get this ?

OpenStudy (anonymous):

no

hartnn (hartnn):

u know how to find volume ?

OpenStudy (anonymous):

v=x^2 h

hartnn (hartnn):

i have written same thing , just denoted height by y instead of h

hartnn (hartnn):

and surface area = ??

OpenStudy (anonymous):

sa= x^2 + 4xh

hartnn (hartnn):

there will be 2 square faces, isn't it ? so 2x^2+4xh ok?

OpenStudy (anonymous):

ok

hartnn (hartnn):

now 2x^2+4xh=337.5 given can u find h in terms of x from here ??

OpenStudy (anonymous):

ahm no

hartnn (hartnn):

subtract 2x^2 from both sides then divide by 4x on both sides. tell me what u get?